OpenStudy (anonymous):

NEED HELP!!!!! Given the function f(x) = 3x + 1, evaluate f(a + 1).

OpenStudy (anonymous):

hi, what you need to do is to put a+1, where there is x

OpenStudy (anonymous):

do you get it

OpenStudy (anonymous):

ok...i will try.

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

look, buddy I have to go, but after doing that you have to remove the bracket the answer is\[f(a+1)=3a+4\], I hope that helps

OpenStudy (anonymous):

it helped.... i tried to sand a massage but my internet was sllloww...thankx

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

how about this one...If f(x) = x 2 + 1, find f(a + 1). should i just (a+1) into x??

OpenStudy (anonymous):

do you know why i said put a+1 in x

OpenStudy (anonymous):

2 find out what x is right????

OpenStudy (anonymous):

not really, the reason is that we were given the funtion of x "(f(x))", than we were asked to avaluate F(a+1), do you see that they just tell us where there is x put x+1

OpenStudy (anonymous):

then how should i do this problem...i have tried 3 times...

OpenStudy (anonymous):

is it like this or not\[f(x)=x^2+1\]

OpenStudy (anonymous):

this is the problem i am working on....If f(x) = x 2 + 1, find f(a + 1).

OpenStudy (anonymous):

you need to do the same thing

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

ok thankx u really helped me a lot..:)

OpenStudy (anonymous):

any thing else

OpenStudy (anonymous):

no, thank you for everything i just handed in my assessment and i had a 86% :)

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

if you need help let me know

OpenStudy (anonymous):

ok i have a another problem....how should i do the problem?? Let R={0, 1, 2, 3} be the range of h(x) = x - 7. The domain of h inverse is